# A Learning-based Domain Decomposition Method

### v1

@@ Abstract
{highlights}
A learning-based domain decomposition method (L-DDM) is proposed with neural operators serving as local surrogate models, efficiently solving parameterized PDEs over large and complicated domains.
The physics-pretrained neural operator (PPNO) is proposed to effectively learn the solution mappings for PDEs from both parameterization and boundary conditions,
@@ 3.4 U-shaped Architecture
where $\omega^{(2T+1)}$ is the output of the neural operator. The operator $\mathcal{F}_{1}:\mathcal{U}(\Omega;\mathbb{R}^{d_{in}})\to\mathcal{U}(\Omega;\mathbb{R}^{d_{1}})$ is called the lifting operator since, usually, $d_{1}$ is chosen so that $d_{1}>d_{in}$. Then each down-sampling block is a mapping $B_{\text{d}}^{(j-1)}:\mathcal{U}(\Omega;\mathbb{R}^{d_{j-1}})\to\mathcal{U}(\Omega;\mathbb{R}^{d_{j}})$ for $2\leq j\leq T$. It is common practice to choose $d_{j}>d_{j-1}$. The middle block is a mapping $M:\mathcal{U}(\Omega;\mathbb{R}^{d_{T}})\to\mathcal{U}(\Omega;\mathbb{R}^{d_{T}})$ and the up-sampling blocks are mappings $B_{\text{u}}^{(j)}:\mathcal{U}(\Omega;\mathbb{R}^{d_{T-j+1}}\times\mathbb{R}^{d_{T-j+1}})\to\mathcal{U}(\Omega;\mathbb{R}^{d_{T-j}})$ for $11\leq j\leq T-$ 1. In particular, they take as input the output of the previous block as well as the down-sampling corresponding of the same latent dimension. This ensures that information is not lost due to the down-sampling operation at the beginning of the architecture. The final operator $\mathcal{F}_{2}:\mathcal{U}(\Omega;\mathbb{R}^{d_{1}})\to\mathcal{U}(\Omega;\mathbb{R}^{d_{out}})$ is usually called the projection as it linearly maps onto the expected number of output channels. A pictorial representation of the architecture is shown in Figure 3.
@@ 4.1 Pretrained PPNO: Learning the Local PDE
Relative $\text{Relative }L_{2}\text{ Error}\text{ (RLS)}=\frac{1}{N}\sum_{n=1}^{N}\frac{\|\hat{u}_{n}-u_{n}\|^{2}_{L^{2}}}{\|u_{n}\|^{2}_{L^{2}}}.$ (45)
A total of 10,000 data samples are generated in the pre-training phase. The models are trained with 8000 samples generated using the Matlab finite element solver [53] with a fixed grid resolution $129\times 129$. 2000 data samples are used to test the trained model. Visualizations of median-error test samples for the prediction of $u$ and $\nabla_{x}{u}$ can be found in Fig. 4. We use Adam optimizer with CosineAnnealing scheduler with a learning rate of 5e-5 and batch size of 10. For a model trained on 10k data for 500 epochs using these hyperparameters and accelerated with an NVIDIA A100 GPU, the training time is approximately 16 hours.
@@ 4.2 Learning-based Domain Decomposition
For the solution process, the global domain is partitioned into a $10\times 10$ grid of identical square subdomains with a 31.25% overlap ratio, as depicted in Figure 7(c). Each subdomain thus contains approximately 50 grains, consistent with the scale of the pre-training data. Random Dirichlet boundary conditions, generated according to Eq. (41) with parameters specified in F, are applied to the exterior edges (Figure 7(c)).
@@
Figure 7: (a) The random Dirichlet boundary conditions applied to the 4 edges of the scaled square domain. (b) The square domain with Voronoi crystal microstructure decomposed into $10\times 10$ subdomains, where the shaded area represents the overlapping of subdomains. (c) 100 subdomains each containing 50 grains after decomposition.
Figure 8: (a-b) The truth solution $u$ and learning-based solution $\hat{u}$ field for the $10\times 10$ rectangular domain, with boundary conditions prescribed in Fig.7c applied. (c) The iterative (black line) and successive (red line) error versus the number of iterations of the $\hat{u}$ field. (d-e) The truth solution $\nabla_{x}u$ and learning-based solution $\nabla_{x}\hat{u}$ field for the same rectangular domain with the same boundary conditions applied. (g) The iterative (black line) and successive (red line) error versus the number of iterations of the $\nabla_{x}\hat{u}$ field.
@@ Acknowledgment
NBK is grateful to the NVIDIA Corporation for support through full-time employment.
@@ Appendix D Rectangular domains with out-of-distribution microstructure patterns
Figure 14: $4\times 3$ rectangular domains with different microstructures, specifically varying grain sizes(a), hexagonal structures(i) and fibre-composite microstructure(q); (e,m,u) The boundary conditions applied; (b,j,r) Truth solution of the $u$ field; (c,k,s) Learning based solution of the $u$ field; (d,l,t) Error versus number of iterations of the $u$ field; (f,n,v) Truth solution of the $\nabla u$ field; (g,o,w) Learning based solution of the $\nabla u$ field; (h,p,x) Error versus number of iterations of the $\nabla u$ field.

### v2

@@ 3.4 U-shaped Architecture
where $\omega^{(2T+1)}$ is the output of the neural operator. The operator $\mathcal{F}_{1}:\mathcal{U}(\Omega;\mathbb{R}^{d_{in}})\to\mathcal{U}(\Omega;\mathbb{R}^{d_{1}})$ is called the lifting operator since, usually, $d_{1}$ is chosen so that $d_{1}>d_{in}$. Then each down-sampling block is a mapping $B_{\text{d}}^{(j-1)}:\mathcal{U}(\Omega;\mathbb{R}^{d_{j-1}})\to\mathcal{U}(\Omega;\mathbb{R}^{d_{j}})$ for $2\leq j\leq T$. It is common practice to choose $d_{j}>d_{j-1}$. The middle block is a mapping $M:\mathcal{U}(\Omega;\mathbb{R}^{d_{T}})\to\mathcal{U}(\Omega;\mathbb{R}^{d_{T}})$ and the up-sampling blocks are mappings $B_{\text{u}}^{(j)}:\mathcal{U}(\Omega;\mathbb{R}^{d_{T-j+1}}\times\mathbb{R}^{d_{T-j+1}})\to\mathcal{U}(\Omega;\mathbb{R}^{d_{T-j}})$ for $1\leq j\leq T-1$. In particular, they take as input the output of the previous block as well as the down-sampling corresponding of the same latent dimension. This ensures that information is not lost due to the down-sampling operation at the beginning of the architecture. The final operator $\mathcal{F}_{2}:\mathcal{U}(\Omega;\mathbb{R}^{d_{1}})\to\mathcal{U}(\Omega;\mathbb{R}^{d_{out}})$ is usually called the projection as it linearly maps onto the expected number of output channels. A pictorial representation of the architecture is shown in Figure 3.
@@ 4.1 Pretrained PPNO: Learning the Local PDE
$\text{Relative }L_{2}\text{ Error}\text{ (RLS)}=\frac{1}{N}\sum_{n=1}^{N}\frac{\|\hat{u}_{n}-u_{n}\|^{2}_{L^{2}}}{\|u_{n}\|^{2}_{L^{2}}}.$ (45)
A total of 10,000 data samples are generated in the pre-training phase. The models are trained with 8000 samples generated using the Matlab finite element solver [53] with a fixed grid resolution $\times 129$ 129. 2000 data samples are used to test the trained model. Visualizations of median-error test samples for the prediction of $u$ and $\nabla_{x}{u}$ can be found in Fig. 4. We use Adam optimizer with CosineAnnealing scheduler with a learning rate of 5e-5 and batch size of 10. For a model trained on 10k data for 500 epochs using these hyperparameters and accelerated with an NVIDIA A100 GPU, the training time is approximately 16 hours.
@@ 4.2 Learning-based Domain Decomposition
For the solution process, the global domain is partitioned into a $\times 10$ 10 grid of identical square subdomains with a 31.25% overlap ratio, as depicted in Figure 7(c). Each subdomain thus contains approximately 50 grains, consistent with the scale of the pre-training data. Random Dirichlet boundary conditions, generated according to Eq. (41) with parameters specified in F, are applied to the exterior edges (Figure 7(c)).
@@
Figure 7: (a) The random Dirichlet boundary conditions applied to the 4 edges of the scaled square domain. (b) The square domain with Voronoi crystal microstructure decomposed into $\times 10$ 10 subdomains, where the shaded area represents the overlapping of subdomains. (c) 100 subdomains each containing 50 grains after decomposition.
Figure 8: (a-b) The truth solution $u$ and learning-based solution $\hat{u}$ field for the $\times 10$ 10 rectangular domain, with boundary conditions prescribed in Fig.7c applied. (c) The iterative (black line) and successive (red line) error versus the number of iterations of the $\hat{u}$ field. (d-e) The truth solution $\nabla_{x}u$ and learning-based solution $\nabla_{x}\hat{u}$ field for the same rectangular domain with the same boundary conditions applied. (g) The iterative (black line) and successive (red line) error versus the number of iterations of the $\nabla_{x}\hat{u}$ field.
@@ Acknowledgment
The work was funded by EPSRC grant EP/W524633/1. The authors are grateful to Dr. Yuxuan Chen for helpful discussions. NBK is grateful to the NVIDIA Corporation for support through full-time employment.
@@ Appendix D Rectangular domains with out-of-distribution microstructure patterns
Figure 14: 4×34$\times$ 34 × 3 rectangular domains with different microstructures, specifically varying grain sizes(a), hexagonal structures(i) and fibre-composite microstructure(q); (e,m,u) The boundary conditions applied; (b,j,r) Truth solution of the $u$ field; (c,k,s) Learning based solution of the $u$ field; (d,l,t) Error versus number of iterations of the $u$ field; (f,n,v) Truth solution of the $\nabla u$ field; (g,o,w) Learning based solution of the $\nabla u$ field; (h,p,x) Error versus number of iterations of the $\nabla u$ field.